The number of candies consumed varies inversely with the number of children present. When 12 children are present, 140 pieces of candy are consumed. Therefore, when 8 children are present, _____ pieces of candy are consumed
Answer by stanbon(75887)

The number of candies consumed varies inversely with the number of children present. When 12 children are present, 140 pieces of candy are consumed. Therefore, when 8 children are present, _____ pieces of candy are consumed
-----
n = k/c
-----
Solve for "k"::
140 = k/12
k = 1680
----
Equation:
n = 1680/c
If c = 8, n = 1680/8 = 210
-------------------
Cheers,
Stan H.
--------------